IF I WANT MAKE A LINE which HAS A SLOPE OF -2 AND PASSES THROUGH (1,4) WHAT I SUPPOSE TO DO?

Answer by richwmiller(17219)

y - y1 = m( x - x1 )
m=slope
y1=your 1
and x1=your 4
plug in those numbers.
y-1=-2(x-4)
and solve for y so that all that is on the left is y
that will be your equation